I have to solve this equation by factoring.
3x^2-6x=24
could you show me how to do this problem?

Answer by waynest(281)

3x^2 -6x = 24
3x^2 - 6x - 24 = 24 - 24
3x^2 - 6x - 24 = 0
(3x + 6)(x - 4) = 0
check:
3x^2 - 12x + 6x - 24 = 0
3x^2 - 6x - 24 = 0
3x^2 - 6x - 24 + 24 = 24
3x^2 - 6x = 24
solve for x:
3x + 6 = 0
3x + 6 - 6 = -6
3x = -6
3x/3 = -6/3
x = -2
x - 4 = 0
x - 4 + 4 = 4
x = 4